A wafer manufacturer monitors the number of defective wafers produced and is developing a control chart to monitor the process. A sample group with 100 wafers in each subgroup was selected and the average percent of defective wafers in that sample was 5%. Determine the LCL of a p chart. Include 2 decimal places in your answer (round (down for x.xx4 and below and up for x.xx5 and up) based on the value of the third decimal place). Remember if the lower control limit is less than 0, its value should be reset to 0 since a proportion cannot have a negative value.
Answer #1

Answer: UCL= 0.1154, LCL= 0

Explanation: Average percent of defects= 5%

Therefore the average proportion of defects= 5/100= 0.05

steps and formulas, for z=3
Proportion of defects=P total defectives/total observations 0.0500
Q= 1-P 0.9500
N= average sample size 100
Standard deviation, Sp squareroot(P*Q/N) 0.0218
UCL= P + z*Sp 0.1154
LCL= P - z*Sp -0.0154
defects cannot be negative, therefore negative LCL is taken as '0' 0.0000